# QOG:Question and Options Generation based on Language Model

### v2

@@ Abstract
Question-Options Generation (QOG) is a task that involves generating a set of question-options pairs given context. This task has various applications, including fine-tuning large models, information retrieval, and automated multiple-choice question generation for education. In this paper, we develop QOG models using three different methods based on fine-tuning sequence-to-sequence language models (LMs). Experiments demonstrate that the end-to-end QOG model is computationally efficient and stable during both training and inference, outperforming other methods. Furthermore, our analysis indicates that our QOG models are competitive on the QOG task compared to the large language model Llama 3-8B.
@@ 1 Introduction
Question and Options Generation (QOG) refers to the task of generating a set of question-options pairs given an input context (e.g. a paragraph). QOG can be used to develop unsupervised question answering and as a data augmentation tool (Alberti et al., 2019; Yu et al., 2018; Riabi et al., 2020; Singh et al., 2019; Longpre et al., 2019) to enhance the text understanding capabilities of large language models. QOG can also be used as an educational aid(Agarwal et al., 2019; Cai et al., 2023; Wang et al., 2023), to improve information retrieval models(Nogueira et al., 2019; Pyatkin et al., 2021), and as a tool for model explanation.
@@
QOG originates from Question Generation (QG)(Duan et al., 2017; Du et al., 2017; Ali et al., 2010), which involves generating a question for an answer given an input context. Compared to QG and QAG, QOG is a more complex task because the correct answer and distractors need to be generated rather than assumed as part of the input. Currently, it is unknown which models are effective for the QOG task because no comprehensive comparative study has been established to date.
@@ 2.1 Question Generation
Question generation (QG) aims to generate questions given documents and answers, which has been applied in various scenarios such as data augmentation(Alberti et al., 2019; Pellicer et al., 2023) and document retrieval(Nogueira et al., 2019; Hofstätter et al., 2020; Ram et al., 2023). Puri et al. (2020) fine-tuned QG’s autoregressive LM to generate related questions. These works focus on generating large-scale data at once without paying much attention to data quality. Xiong and Wu (2020) followed the idea of adversarial generation and used question-answering agents and QG models to train adversarial models. This work used semantic similarity as an evaluation indicator to improve the quality of generated data, but did not focus on the generalization ability of the model in other fields. To this end, our work will focus on improving data quality while testing the generalization ability of the model with data sets from different fields.
@@ 4.1 Experimental Setting
Dataset Size
SQuAD-QO 87,399
SQuADShifts-QO 37,691
FinQA-QO 8,281
Table 1: Size of Datasets

### v3

@@ Abstract
Question-Options Generation (QOG) is a task that involves generating a set of question-options pairs within a given input. This task has various applications, including fine-tuning large models, information retrieval, and automated multiple-choice question generation for education. In this paper, we develop QOG models using three different methods based on fine-tuning sequence-to-sequence language models (LMs). Experiments demonstrate that the end-to-end QOG model is computationally efficient and stable during both training and inference, outperforming other methods. Furthermore, our analysis indicates that our QOG models are competitive on the QOG task compared to the large language model Llama 3-8B.
@@ 1 Introduction
Question and Options Generation (QOG) refers to the task of generating a set of question-options pairs given an input context (e.g. a paragraph). QOG can be used to develop unsupervised question answering and as a data augmentation tool (Alberti et al., 2019; Yu et al., 2018; Riabi et al., 2020; Singh et al., 2019; Longpre et al., 2019) to enhance the text understanding capabilities of large language models. QOG can also be used as an educational aid(Agarwal et al., 2019; Cai et al., 2023; Wang et al., 2023), to enhance information retrieval models(Nogueira et al., 2019; Pyatkin et al., 2021), and as a way to explain models.
@@
QOG originates from Question Generation (QG)(Duan et al., 2017; Du et al., 2017; Ali et al., 2010), which involves generating a question for an answer given an input context. Compared to QG and QAG, QOG presents a more intricate challenge, as the correct answer and distractors must be constructed rather than assumed as part of the input. Currently, it is unknown which models are effective for the QOG task because no comprehensive comparative study has been established to date.
@@ 2.1 Question Generation
Question generation (QG) aims to generate questions given documents and answers. This approach has been applied in a variety of scenarios, including data augmentation(Alberti et al., 2019; Pellicer et al., 2023) and document retrieval(Nogueira et al., 2019; Hofstätter et al., 2020; Ram et al., 2023). Puri et al. (2020) fine-tuned QG’s autoregressive LM to generate related questions. These works focus on generating large-scale data at once without paying much attention to data quality. Xiong and Wu (2020) followed the idea of adversarial generation and used question-answering agents and QG models to train adversarial models. This work used semantic similarity as an evaluation indicator to improve the quality of generated data, but did not focus on the generalization ability of the model in other fields. To this end, our work will focus on improving data quality while testing the generalization ability of the model with data sets from different fields.
@@ 4.1 Experimental Setting
Dataset Size
SQuAD-QO 87,399
SQuADShifts-QO 37,691
FinQA-QO 8,281
Table 2: Size of Datasets